# Slow magnetic quantum oscillations in the $c$-axis magnetoresistance of UTe2

### v1

@@ III Results and Discussion
where $A_{i}$ are oscillation amplitudes for each frequency, $R_{i,\mathrm{D}}=\exp{(\alpha m_{i}^{*}T_{\mathrm{D}}/B)}$ are Dingle damping factors, $R_{i,\mathrm{MB}}$ are MB damping factors, $R_{\mathrm{i,S}}$ are spin damping factors, and $\gamma_{i}$ oscillation phases. In Fig. 5, we demonstrate this by a simple simulation using a superposition of the two frequencies, $F_{1}=85\,$T and $F_{2}=205\,$T, shown in the upper panel as green and blue curves. We introduce the damping factors $R_{1,\mathrm{MB}}=\left[1-\exp{\left(-B/B_{\mathrm{MB}}\right)}\right]^{n/2}$ and $R_{2,\mathrm{MB}}=\exp{\left(-B/B_{\mathrm{MB}}\right)}^{n/2}$ with the number of breakdown junctions $n$, which we assume to be 2 (see below). A small breakdown field of $B_{\mathrm{MB}}=6\,$T is sufficient to yield a curve that is roughly similar to the experimental one (magenta curve in Fig. 5 lower panel). Although the MB effect may account for the emergence of $F_{2}$, it cannot explain the sudden suppression of $F_{1}$ as observed. Our results indicate that $F_{1}$ vanishes very abruptly around 16 T and remains completely suppressed for higher fields. In a conventional MB scenario, the two frequencies are expected to transition rather smoothly as explained above.
@@
Let us now look at the FS of UTe2. In Figs. 6(a) and 6(b) we show results from DFT calculations for two different $U$ values of 1 and 2 eV. The latter reproduces previously reported results [15]. For smaller $U$, the electron band exhibits a more 3D character with a closed loop around the $\Gamma$ point in the first Brillouin zone [marked by the yellow dashed line in Fig. 6(a)]. Hence, such a closed trajectory could account for a MQO observable for $H\parallel a$. Nevertheless, the result shown in Fig. 6(b), exhibiting two quasi-2D FS cylinders, has been considered more realistic, since it is in line with recent dHvA measurements [41, 42]. For the predicted quasi-2D FSs oriented along the $c$ direction, an external field applied along $a$ would force the charge carriers along open trajectories [following the vertical yellow dashed lines in Fig. 6(b)], rendering closed cyclotron orbits impossible. As these bands possess effective masses between 32 and 57 free-electron masses [41], we would expect a similar mass value for a conventional MB orbit. The mass should even be further enhanced as the field is tilted towards the in-plane direction, which is clearly not the case neither for $F_{1}$ nor for $F_{2}$. Moreover, if contributions from both electron- and hole-like FSs were involved, the small frequency values may only be possible via classically forbidden trajectories, as the Lorentz force would switch its sign as soon as the charge carriers move from one to the other band via a potential breakdown gap. Hence, the observed MQOs for $H\parallel a$ are rather unexpected and likely of an unconventional origin.
@@
Alternative origins such as QIOs, according to a proposal by Eaton et al. [42], oscillations due to inter- and intraband scattering (i.e. QPLOs) [48], oscillations due to weakly warped 2D FSs [52, 53], or oscillations of the chemical potential [50] may account for the observed small frequencies. Oscillations in the chemical potential require large oscillation amplitudes due to strongly 2D FSs. Also, these MQOs should be picked up by the magnetization. By contrast, both QIOs and QPLOs should be most pronounced in electrical-transport probes [48]. Our measurements of magnetic torque on a single crystal (same growth batch of sample S2) and also those of another group [41], to date, have only revealed oscillations associated with the quasi-2D FS cylinders. The slow oscillations appear to be undetectable by this thermodynamic probes, rendering oscillations in the chemical potential an unlikely origin for $F_{1}$ and $F_{2}$. For weakly warped quasi-2D FSs, slow oscillations can emerge that possess extremely low effective masses, as it was reported for organic superconductors [52, 53]. This effect, however, strongly relies on field orientations aligned along the quasi 2D FS cylinders, which is counter to the observations in this work (strongest MQOs for $I\parallel c$, but with $H\parallel a$).
@@
Recent transport studies with current along the $a$ direction confirmed the heavy FSs but did not resolve these slow MQOs either [58]. We, therefore, tested an additional sample S3 cut from the same crystal used for S2, however, with current aligned along $a$. No low-frequency MQOs were discernible for $H\parallel a$ under the same noise conditions at $T=0.6\,$K and field up to 16 T. This implies that it is the $c$-axis transport that is most sensitive to the low-frequency MQOs.
@@
Distinguishing QIOs from QPLOs is challenging [48]. In Fig. 6(c) and 6(d), we show the 2D projection of the two distinct FS cylinders along the $k_{a}$ direction. A small area is discernible where the projections of the two bands do not overlap completely, providing a narrow region where two pathways could lead to MQOs based on quantum interference via the narrow gaps in between [44]. QIOs are associated with MB via a narrow gap, $\Delta_{\mathrm{MB}}$, between two different parts of the FS [Fig. 6(e)] and require a finite tunneling probability. At large enough magnetic field, the oscillation amplitude is expected to subside again as the breakdown probability reaches $100\,\%$ Hence, only one path will be allowed for in the interference pattern sketched in Fig. 6(e).
@@
In contrast, QPLOs originate from inter- and intraband scattering processes and would exhibit ever growing amplitudes with increasing field [59]. In our high-field data, we do not observe a significant damping of the oscillation strength up to 60 T, that is, for $F_{2}$ the oscillation amplitude increases monotonically upon increasing field up to 60 T [Fig. 3(b)]. Moreover, QPLOs would be very stable against temperature and can enable closed orbits between distant parts of a FS, rendering them a likely alternative origin for the observed MQOs. An additional point against the suggested quantum interference is that the orbits, required for MQOs when $H\parallel a$, run along the heavy open quasi-2D bands and are connected via MB gaps only at specific touching points. Therefore, very few charge carriers would be expected to contribute to the $c$-axis magnetotransport, which would result in an overall negligible contribution to the main conduction channel.
@@
In Fig. 6(d), we present DFT results with an artificial variation in the position of the chemical potential by ±$\pm 50\,$meV. This demonstrates that the overlap between the projections of the two bands strongly depends on the details of the electronic structure, which certainly are beyond the reach of the present state of the art of DFT calculations in UTe2. Note: Other calculations based on dynamic mean field theory show significant differences to the DFT results [25, 60]. As we observe a rather abrupt transition from $F_{1}$ to $F_{2}$ induced by field, the question arises how magnetic field may affect the FS properties and what mechanism can account for the observed transition. Even though QIOs or QPLOs seem plausible, at this point, we cannot rule out a more complex FS topology that may even include a smaller 3D FS pocket responsible for the low-frequency MQOs $F_{1}$ and $F_{2}$.
Another argument for a 3D FS is the rather small anisotropy of the electrical transport in UTe2 [22, 61]. Moreover, the overall MR for $H\parallel a$ appears to be rather small, e.g., at 0.5 K it changes by a factor of two between 10 and 24 T [Fig. 1(a)]. This is surprisingly weak for a quasi-2D FS with open bands along the $c$ axis. Such open bands should result in a non-saturating MR for $H\parallel a$ and a strongly anisotropic MR between this orientation and $H\parallel c$. Even though for $H\parallel a$, the MR exhibits no sign of saturation, we observe an equal or even smaller MR than for the other two principle axes.
@@
Figure 7: (a) $c$-axis MR recorded for various fixed angles $\phi$ with 20 µA excitation current at 222\,2K. (b) Hall resistance recorded simultaneously with the data shown in (a). Red dots mark zero crossings. (c) Residual MR for $T=0.5\,$K after subtracting the slowly varying background fitted by a polynomial of $4^{\mathrm{th}}$ degree. Red dots on the flat part before the oscillations emerge mark the values where the Hall resistance, shown in (b), changes its sign. The initial upturn originates from the superconducting transition at low fields. Black arrows highlight the first hump-like features. Blue arrows highlight the first minima.
@@
In Figs. 7(a)-7(c), we compare MR data obtained for sample S2 for various field orientations within the $a$-$b$ plane at 2 K with simultaneously measured Hall-effect data and the resolved MQOs superimposed to the MR at 0.5 K. There is a broad hump in the normal-state MR for $H\parallel a$ around $6-7\,$T [Fig. 7(a)] that shifts towards higher fields upon rotating away from $a$. The field value of this feature is much lower than the one at which the MQOs oscillations set in [marked by black and blue arrows in Fig. 7(c)]. The Hall resistance exhibits a maximum and clear sign change (marked by red dots) that both shift to higher fields upon increasing tilt angles $\phi$. Previous thermoelectric measurements with $I\parallel a$ and $H\parallel a$ observed features around 6 T in various properties such as the Seebeck coefficient, the Hall effect, and the MR [33]. A transitional change in the FS properties, associated with a Lifshitz transition at this field value, was proposed. In general, such a transition is connected with a topological change of the FS, i.e., a change in the charge-carrier distribution at the Fermi edge [62, 63]. Here, the question arises, what effect such a transition has on the FS in UTe2. The slow oscillations only start emerging at about 11 T, just above the field range where the Hall effect changes its sign from positive to negative [marked by red circles in Figs. 7(b) and 7(c)]. This implies that the FS undergoes significant changes upon increasing field. Interestingly, Seebeck measurements revealed multiple features at fields of about 6, 10, and 20 T [33]. Similarly, we observe $F_{1}$ and $F_{2}$ to emerge consecutively at different fields in our experiments.
@@
Our results would, therefore, support the presence of at least two field-induced Lifshitz transitions associated with the emergence of $F_{1}$ and $F_{2}$. A recent preprint highlights the expected effect of coherent and incoherent Kondo scattering on the FS, causing temperature-, field-, and pressure-induced topological FS transformations [64]. This proposal could also account for the observed MQOs that emerge and vanish at specific magnetic field values or orientations.
@@ IV Conclusion
In summary, we were able to detect slow magnetic quantum oscillations in the resistance of bulk and microfabricated UTe2 samples in different experimental setups with samples from different growths. Two distinct slow frequencies, $F_{1}\approx 90\,$T and $F_{2}\approx 205\,$T, for field along the crystallographic $a$ direction, are discernible, that emerge consecutively upon increasing the external magnetic field. $F_{1}$ starts emerging at around 11 T, the field range at which the Hall effect crosses zero and changes its sign. At around 16 T, $F_{1}$ ceases to exist and instead a new frequency $F_{2}$ sets in.
We determine effective masses of $1.5m_{\mathrm{e}}$ for both frequencies and trace them depending on tilt angles in the $a$-$c$ and $a$-$b$ plane. The oscillations are discernible within only a narrow angular region around the $a$ axis. Although $F_{2}$ matches previous reports from contactless surface-conductivity measurements [44], the lower frequency $F_{1}$ has not been observed to date. We argue that these oscillations likely originate from non-Onsager mechanisms such as quantum interference or quasiparticle lifetime oscillations.
Our observations support the presence at least two field-induced Lifshitz transitions that alter the FS topology of UTe2. Apparently, UTe2 undergoes significant changes induced by external magnetic fields both along the $a$ and $b$ axis. This highlights again the importance of the field as well as its orientation for the complex magnetic phase diagram of UTe2. In order to understand its implications on the superconducting ground state, further evidence for a field-induced Fermi-surface reconstruction and its exact details would be highly desirable.

### v2

@@ III Results and Discussion
where $A_{i}$ are oscillation amplitudes for each frequency, $R_{i,\mathrm{D}}=\exp{(\alpha m_{i}^{*}T_{\mathrm{D}}/B)}$ are Dingle damping factors, $R_{i,\mathrm{MB}}$ are MB damping factors, $R_{\mathrm{i,S}}$ are spin damping factors, and $\gamma_{i}$ oscillation phases. In Fig. 5, we demonstrate this by a simple simulation using a superposition of the two frequencies, $F_{1}=85\,$T and $F_{2}=205\,$T, shown in the upper panel as blue and red curves. We introduce the damping factors $R_{1,\mathrm{MB}}=\left[1-\exp{\left(-B/B_{\mathrm{MB}}\right)}\right]^{n/2}$ and $R_{2,\mathrm{MB}}=\exp{\left(-B/B_{\mathrm{MB}}\right)}^{n/2}$ with the number of breakdown junctions $n$, which we assume to be 2 (see below). A small breakdown field of $B_{\mathrm{MB}}=6\,$T is sufficient to yield a curve that is roughly similar to the experimental one (magenta curve in Fig. 5 lower panel). Although the MB effect may account for the emergence of $F_{2}$, it cannot explain the sudden suppression of $F_{1}$ as observed. Our results indicate that $F_{1}$ vanishes very abruptly around 16 T and remains completely suppressed for higher fields. In a conventional MB scenario, the two frequencies are expected to transition smoothly as explained above.
@@
Let us now look at the FS of UTe2. In Figs. 6(a) and 6(b), we show results from DFT calculations for two different $U$ values of 1 and 2 eV. We chose the latter value to reproduce previously reported results [15] and the former to showcase the development of a 3D pocket around the center of the Brillouin zone. For smaller $U$, the electron band exhibits a more 3D character with a closed loop around the $\Gamma$ point in the first Brillouin zone [marked by the yellow dashed line in Fig. 6(a)]. Hence, such a closed trajectory could account for a MQO observable for $H\parallel a$. Nevertheless, the result shown in Fig. 6(b), exhibiting two quasi-2D FS cylinders, is considered more realistic, since it is in line with recent dHvA measurements [41, 42]. For the predicted quasi-2D FSs oriented along the $c$ direction, an external field applied along $a$ would force the charge carriers along open trajectories [following the vertical yellow dashed lines in Fig. 6(b)], rendering closed cyclotron orbits impossible. As these bands possess effective masses between 32 and 57 free-electron masses [41], we would expect a similar mass value for a conventional MB orbit. The mass should even be further enhanced as the field is tilted towards the in-plane direction, which is clearly not the case neither for $F_{1}$ nor for $F_{2}$. Moreover, if contributions from both electron- and hole-like FSs are involved, the small frequency values may only be possible via classically forbidden trajectories, as the Lorentz force would switch its sign as soon as the charge carriers moved from one to the other band via a potential breakdown gap. Hence, the observed MQOs for $H\parallel a$ are rather unexpected and likely of an unconventional origin.
Various mechanisms may account for the observed low frequencies: a weak warping of quasi-2D FSs [49, 50]; oscillations of the chemical potential  [46]; inter- and intraband scattering (i.e., quasi-particle lifetime oscillations) [48]; or quantum interference, according to a proposal for UTe2 by Eaton et al. [42].
For weakly warped quasi-2D FSs, slow oscillations can emerge that possess extremely low effective masses, as it was reported for organic superconductors [49, 50]. This effect, however, strongly relies on field orientations aligned along the quasi-2D FS cylinders, which is counter to the observations in this work (strongest MQOs for $I\parallel c$, but with $H\parallel a$).
@@
Oscillations in the chemical potential can be ruled out in our case, as they usually occur with large oscillation amplitudes due to strongly 2D FSs. Also, these MQOs should be picked up by the magnetization, which, so far, has not been the case here. Our measurements of magnetic torque on a single crystal (same growth batch of sample S2) and also those of another group [41], to date, have only revealed oscillations associated with the quasi-2D FS cylinders. The slow oscillations appear to be undetectable by these thermodynamic probes, rendering oscillations in the chemical potential an unlikely origin for $F_{1}$ and $F_{2}$. By contrast, oscillations that originate from ntra- or interband scattering or from quantum interference should be most pronounced in electrical-transport probes [48].
In the first case, quasiparticles may scatter on defects or collective excitations, connecting different parts of the FS. The corresponding additional MQOs have effective masses that result from a combination of the underlying bands [47]. In the second case of quantum interference, MQOs arise from coherent charge-carrier pathways between different MB junctions, effectively enclosing flux in an interferometer-like configuration [57, 53, 48]. MB is a well-known effect in the magnetic field based on tunnel transitions of conduction electrons of a metal between trajectories belonging to different electronic bands [54, 1]. According to the authors of Ref. [54], the observation of quantum interference requires a FS that (i) realizes an interferometer setup, i.e., at least two semiclassical paths on which particles travel from one MB junction to the next and which enclose an area in momentum space; (ii) MB junctions between the semiclassical paths with intermediate MB probabilities $p$ leading to a partial MB with $p(1-p)\neq 0$. One of the key signatures of quantum interference oscillations (QIOs) is their stability against increased temperature [1].
According to theory, the effective mass of QIOs results from the difference of the two involved parts of the FS. Hence, it can be significantly smaller than compared to the fundamental renormalized band masses, reducing the damping of the oscillation amplitude due to temperature. Specifically, this feature may enable the detection of the MQOs at higher temperatures in the case of the heavy fermion UTe2 [42].
@@
Recent transport studies with current along the $a$ direction confirmed the heavy FSs, but did not resolve these slow MQOs either, at least for a field applied in the $b$-$c$ plane [58]. We, therefore, tested an additional sample S3 cut from the same crystal used for S2, however, with the current aligned along $a$. No low-frequency MQOs were discernible for $H\parallel a$ under the same noise conditions at $T=0.6\,$K and field up to 16 T. This implies that it is the $c$-axis transport that is most sensitive to the low-frequency MQOs.
@@
Distinguishing quantum interference from intra-/interband scattering is challenging [48]. In Figs. 6(c) and 6(d), we show the 2D projection of the two distinct FS cylinders along the $k_{a}$ direction. A small area is discernible where the projections of the two bands do not overlap completely, providing a narrow region where two pathways could lead to MQOs based on quantum interference via the narrow gaps in between [44]. QIOs are associated with MB via a narrow gap, $\Delta_{\mathrm{MB}}$, between two different parts of the FS [Fig. 6(e)] and require a finite tunneling probability. At large-enough magnetic field, the oscillation amplitude is expected to subside again as the breakdown probability reaches $100\,\%$. Hence, only one path will be allowed for in the interference pattern sketched in Fig. 6(e).
@@
In contrast, the so-called quasi-particle lifetime oscillations (QPLOs) originate from intra- and interband scattering processes and would exhibit ever-growing amplitudes with increasing field [59]. In our high-field data, we do not observe a significant damping of the oscillation strength up to 60 T, that is, for $F_{2}$ the oscillation amplitude increases monotonically upon increasing field up to 60 T [Fig. 3(b)]. Moreover, QPLOs would be very stable against temperature and can enable closed orbits between distant parts of a FS, rendering them a likely alternative origin for the observed MQOs. An additional point against the suggested quantum interference is that the orbits, required for MQOs when $H\parallel a$, run along the heavy open quasi-2D bands and are connected via MB gaps only at specific touching points. Therefore, very few charge carriers would be expected to contribute to the $c$-axis magnetotransport, which would result in an overall negligible contribution to the main conduction channel. However, at this point, making a clear distinction between QIOs and QPLOs goes beyond the scope of this work.
@@
In Fig. 6(d), we present DFT results with an artificial variation in the position of the chemical potential by ±$\pm 50\,$meV. This demonstrates that the overlap between the projections of the two bands strongly depends on the details of the electronic structure, which certainly are beyond the reach of the present state of DFT calculations in UTe2. Note that other calculations based on dynamic mean-field theory show significant differences to the DFT results [25, 60]. As we observe an abrupt transition from $F_{1}$ to $F_{2}$ induced by field, the question arises how magnetic field may affect the FS properties and what mechanism can account for the observed transition. Even though QIOs or QPLOs seem plausible, at this point, we cannot rule out a more complex FS topology that may even include a smaller 3D FS pocket responsible for the low-frequency MQOs $F_{1}$ and $F_{2}$.
@@
Figure 7: (a) $c$-axis MR recorded for various fixed angles $\phi$ with 20 µA excitation current at 222\,2K. (b) Hall resistance recorded simultaneously with the data shown in (a). Red dots mark zero crossings. (c) Residual MR for $T=0.5\,$K after subtracting the slowly varying background fitted by a polynomial of fourth degree. Red dots on the flat part before the oscillations emerge mark the values where the Hall resistance, shown in (b), changes its sign. The initial upturn originates from the superconducting transition at low fields. Black arrows highlight the first hump-like features. Blue arrows highlight the first minima.
The presence of a 3D FS is supported by the small anisotropy of the electrical transport in UTe2 [22, 61]. Moreover, the overall MR for $H\parallel a$ appears to be relatively small, e.g., at 0.5 K it changes by a factor of 2 between 10 and 24 T [Fig. 1(a)]. This is surprisingly weak for a quasi-2D FS with open bands along the $c$ axis. Such open bands should result in a non-saturating MR for $H\parallel a$ and a strongly anisotropic MR between this orientation and $H\parallel c$. Even though for $H\parallel a$ the MR exhibits no sign of saturation, we observe an equal or even smaller MR than for the other two principle axes.
$F_{2}$ is very similar to the MQO previously reported by Weinberger et al. [44]. Even the angular dependence agrees rather well. For rotations from $a$ towards $c$, we resolve oscillations up to about $20^{\circ}$ tilt, matching almost perfectly with the previous report. In the case of rotation from $a$ towards $b$, we are able to trace the oscillation up to $40^{\circ}$. In Ref. [44], additional data for rotation within the $a$-$b$ plane is provided. There, a faint peak is still resolvable in the FFT spectra for tilts larger than $20^{\circ}$ up to $28^{\circ}$. Even though this is a lower angle as compared to our result, it also reveals a slight anisotropic behavior in the angle dependence of the $\sim 200\,$T frequency oscillation. Hence, the limited angular range for $F_{2}$ disagrees with the finding in Ref. [40], where a $\sim 200\,$T frequency was discernible over the full angular range within both the $a$-$b$ and the $a$-$c$ plane, which was associated with a 3D FS pocket. The apparent anisotropic angle dependence in our case may signal a significant anisotropy in the FS properties. Our findings, therefore, support the proposed quantum-interference scenario, with interfering trajectories stretched along the $c$ direction [shown in Fig. 6(e)].
@@
In Figs. 7(a) to 7(c), we compare MR data obtained for sample S2 for various field orientations within the $a$-$b$ plane at 2 K with simultaneously measured Hall-effect data and the resolved MQOs superimposed to the MR at 0.5 K. There is a broad hump in the normal-state MR for $H\parallel a$ around 6-7 T [Fig. 7(a)] that shifts towards higher fields upon rotating away from $a$. The field value of this feature is much lower than the one at which the MQOs oscillations set in [marked by black and blue arrows in Fig. 7(c)]. The Hall resistance exhibits a maximum and clear sign change (marked by red dots) that both shift to higher fields upon increasing tilt angles $\phi$. In general, the Hall effect can be comprised of orbital, anomalous, and topological contributions depending on the material [62]. A sign change in the Hall resistivity may have various origins, such as changes in the electronic or magnetic structure. Previous thermoelectric measurements with $I\parallel a$ and $H\parallel a$ observed features around 6 T in various properties such as the Seebeck coefficient, the Hall effect, and the MR [33]. A transitional change in the FS properties, associated with a Lifshitz transition at this field value, is proposed. In general, such a transition is connected with a topological change of the FS, i.e., a change in the charge-carrier distribution at the Fermi edge [63, 64]. Here, the question arises of what effect such a transition has on the FS in UTe2. The slow oscillations only start emerging at about 11 T, just above the field range where the Hall effect changes its sign from positive to negative [marked by red circles in Figs. 7(b) and 7(c)]. This implies that the FS undergoes significant changes upon increasing field. Interestingly, Seebeck measurements reveal multiple features at fields of about 6, 10, and 20 T [33]. Similarly, we observe $F_{1}$ and $F_{2}$ to emerge consecutively at different fields in our experiments.
@@
Our results would, therefore, support the presence of at least two field-induced Lifshitz transitions associated with the emergence of $F_{1}$ and $F_{2}$. As we observe differences in the MQOs between the microstructured and bulk sample, pressure induced by the substrate of the epoxy encapsulation may affect the FS of UTe2. For example, the superconducting ground state in strongly correlated systems may be controllably affected by built-in pressure using microstructures [65]. A pressure-dependent study demonstrated pressure-induced changes in the slow MQOs associated with quantum interference [66]. Furthermore, a theoretical study highlighted the expected effect of coherent and incoherent Kondo scattering on the FS, causing temperature-, field-, and pressure-induced topological FS transformations [67]. This proposal could also account for the observed MQOs that emerge and vanish at specific magnetic field values or orientations.
@@ IV Conclusion
In summary, we were able to detect slow magnetic quantum oscillations in the resistance of bulk and microfabricated UTe2 samples in different experimental setups with samples from different growths. Two distinct slow frequencies, $F_{1}\approx 90\,$T and $F_{2}\approx 205\,$T, for a field along the crystallographic $a$ direction, were discernible, that emerged consecutively upon increasing the external magnetic field. $F_{1}$ started emerging at around 11 T, the field range at which the Hall effect crosses zero and changes its sign. At around 16 T, $F_{1}$ ceased to exist and instead a new frequency $F_{2}$ sets in.
We determined the effective masses of $1.5m_{\mathrm{e}}$ for both frequencies and traced them depending on tilt angles in the $a$-$c$ and $a$-$b$ plane. The oscillations were discernible within only a narrow angular region around the $a$ axis. Although $F_{2}$ matched previous reports from contactless surface-conductivity measurements [44], the lower frequency $F_{1}$ has not been observed to date. We argue that these oscillations likely originated from non-Onsager mechanisms such as quantum interference or quasiparticle lifetime oscillations.
Our observations support the presence of at least two field-induced Lifshitz transitions that alter the FS topology of UTe2. Apparently, UTe2 undergoes significant changes induced by external magnetic fields both along the $a$ and $b$ axes. This highlights again the importance of the field as well as its orientation for the complex magnetic phase diagram of UTe2. To understand its implications on the superconducting ground state, further evidence for a field-induced Fermi-surface reconstruction and its exact details would be highly desirable.